# From Generation to Detection: A Multimodal Multi-Task Dataset for Benchmarking Health Misinformation

### v1

@@ Abstract
Infodemics and health misinformation have significant negative impact on individuals and society, exacerbating confusion and increasing hesitancy in adopting recommended health measures. Recent advancements in generative AI, capable of producing realistic, human-like text and images, have significantly accelerated the spread and expanded the reach of health misinformation, resulting in an alarming surge in its dissemination. To combat the infodemics, most existing work has focused on developing misinformation datasets from social media and fact-checking platforms, but has faced limitations in topical coverage, inclusion of AI-generation, and accessibility of raw content. To address these issues, we present MM-Health, a large scale multimodal misinformation dataset in the health domain consisting of 34,746 news article encompassing both textual and visual information. MM-Health includes human-generated multimodal information (5,776 articles) and AI-generated multimodal information (28,880 articles) from various SOTA generative AI models. Additionally, We benchmarked our dataset against three tasks—reliability checks, originality checks, and fine-grained AI detection—demonstrating that existing SOTA models struggle to accurately distinguish the reliability and origin of information. Our dataset aims to support the development of misinformation detection across various health scenarios, facilitating the detection of human and machine-generated content at multimodal levels.
From Generation to Detection: A Multimodal Multi-Task Dataset for Benchmarking Health Misinformation
Zhihao Zhang1thanks: Equal contribution., Yiran Zhang [^1] footnotemark: 1, Xiyue Zhou2, Liting Huang3, Imran Razzak4, Preslav Nakov4, Usman Naseem1thanks: Corresponding author: usman.naseem@mq.edu.au, 1Macquarie University, Australia, 2University of Sydney, Australia 3University of Technology Sydney, Australia, 4MBZUAI, UAE
@@ 1 Introduction
Health misinformation refers to information that is inaccurate, misleading, or false according to the best available health evidence at the time. Such misinformation tends to spread at unprecedented speed and scale on the World Wide Web and social media Office of the Surgeon General (2021) (OSG), and has been shown to have significant negative effects on society Borges do Nascimento et al. (2022); Abbott et al. (2021); Wang et al. (2019); Muhammed T and Mathew (2022); Nathan Walter and Suresh (2021). During crises, such as outbreaks of infectious diseases like COVID-19, the overproduction of health misinformation in both digital and physical environments is defined as an infodemic. As society transitions into a long COVID era Davis et al. (2023), the infodemic continues to evolve, leading to a reduction in public trust in health professionals Nathan Walter and Suresh (2021). Furthermore, unfiltered exposure to health misinformation can delay or prevent effective disease treatment and even threaten the lives of individuals Wang et al. (2019). Additionally, the recent advent of generative image models, such as Stable Diffusion Rombach et al. (2022), MidjourneyV5 Holz et al. , DALL-E2 Ramesh et al. (2022), as well as human-level text generators like ChatGPT Ouyang et al. (2022) and LLaMa Touvron et al. (2023), has amplified both the quantity and spread of misinformation Zhou et al. (2023a). The ease with which generative AI models can replicate or manipulate multimodal content, including text and images Huang et al. (2024), further contributes to the emergence of health misinformation Park (2024). This presents a significant new challenge in combating the infodemic.
Name Year Human Context Machine Context Multiple AI Models Reliability Originality General Avaliable
@@
Recent studies have focused on developing and analyzing datasets and detection methods to combat the infodemic. Common methods for collecting multimodal health misinformation include scraping social media platforms Kinsora et al. (2017); Hayawi et al. (2022); Cui and Lee (2020); Zhou et al. (2020); Srba et al. (2022); Sun et al. (2023); Li et al. (2020); Chen et al. (2021b), such as Twitter (now known as X), and parsing fact-checking websites Li et al. (2020); Chen et al. (2021b); Srba et al. (2022); Sun et al. (2023), such as Snopes and Poynter. However, these datasets exhibit the following notable limitations:
  • Exclusion of AI-generated misinformation: The majority of existing datasets focus solely on human-generated misinformation, ignoring the growing trend of AI-generated misinformation Zhou et al. (2023b); Bashardoust et al. (2024); Xu et al. (2023). With AI models now capable of generating and manipulating both text and image content Liz-López et al. (2024), it is crucial to include AI-generated misinformation to ensure data diversity.
@@
The limitation of the previous health misinformation datasets are summarised in Table 1. Most of the existing methods for automatically detecting health misinformation include content-based CNNs Cui and Lee (2020); Zhou et al. (2020); Dai et al. (2020), attention-based hierarchical encoders Kinkead et al. (2020), and graph-based attention methods Cui et al. (2020). These methods have demonstrated competitive performance after being trained on health misinformation benchmarks. However, they require extensive training on benchmark datasets to achieve expected performance, which limits the models’ ability to generalise to new health misinformation. Since health misinformation is constantly evolving Okoro et al. (2024), it is important to evaluate generalised models under zero-shot settings. Additionally, these models only accept structured input and provide labelled output without human-readable explanations. This constraint limits the usability of existing detection methods for the general public outside of the research community. To address the aforementioned limitations, we develop MM-Health, a comprehensive multimodal dataset designed for detecting both human and AI generated health misinformation. Additionally, we conduct extensive experiments using state-of-the-art (SOTA) Vision-Language Models (VLLMs) to evaluate their performance in assessing both the reliability and originality of the health information in MM-Health, as well as performing a fine-grained AI detection analysis. Our key contributions are as follows:
  • We create and release a new large-scale multimodal (text and images) dataset called MM-Health, which contains 34,746 news articles. These articles were collected from existing multimodal health datasets and generated using current state-of-the-art (SOTA) AI generative models.
@@ 2 Related Work
Web and social media provide valuable sources of information and play important roles in multiple tasks like depression and suicide identification Naseem et al. (2024), health mention classification Naseem et al. (2022) and creditable knowledge management Nisar et al. (2019). However, systematic reviews shows that health-related misinformation on web and social media posts critical threat to the community, which could lead to delay or prevent treatment and even threaten the lives of individuals Wang et al. (2019).
@@ 2.1 Existing Datasets
MedHelp Kinsora et al. (2017), COAID Cui and Lee (2020), and ANTi-Vax Hayawi et al. (2022) are text datasets designed to address online health misinformation. Specifically, MedHelp is collected from an online general health discussion forum. It contains 1,338 non-misinformation comments and 887 misinformation comments annotated by human experts between 2001 and 2013. COAID includes both news articles and Twitter content focused on COVID-19 between December 2019 and September 2020. It labels 204 fake news articles, 3,565 true news articles, 28 fake claims, and 454 true claims from a total of 1,896 news articles, 516 Twitter posts, and 183,569 Twitter engagements. ANTi-Vax is built solely from Twitter content focused on the COVID-19 vaccine between December 2020 and January 2021. It includes a total of 15,073 tweets, 5,751 of which are misinformation and 9,322 are general vaccine-related tweets.
Figure 1: Data collection process for MM-Health includes: 1) utilising multiple existing open-source health misinformation datasets as data sources, 2) validating the available data samples and collecting human-generated multimodal data from the provided URLs, and 3) implementing generative AI models to collect AI-generated replicated multimodal data. To ensure data quality, both human and AI generated content are evaluated by five human evaluators proficient in English.
MM-COVID Li et al. (2020), ReCOVery Zhou et al. (2020), Monant Srba et al. (2022), MMCoVaR Chen et al. (2021b), and Med-MMHL Sun et al. (2023) are multimodal datasets for health-related misinformation detection. MM-COVID contains 3,981 fake articles and 7,584 real articles related to COVID-19 which collected from Feb 2020 to Jul 2020. ReCOVery includes both news articles and tweets mentioning COVID-19, labelled by NewsGuard  [^2] https://www.newsguardtech.com/ and Media Bias/Fact Check  [^3] https://mediabiasfactcheck.com/. It comprises 118,339 reliable articles and 28,274 unreliable articles sourced between January 2020 and May 2020. MMCoVaR focuses on vaccine-related misinformation between February 2020 and March 2021, featuring 958 unreliable and 1,635 reliable news articles, plus 24,184 tweets categorized as reliable, unreliable, or inconclusive. Med-MMHL covers multiple diseases, integrating LLM-generated fake news, with 1,979 real and 1,604 fake articles, alongside 1,334 reliable and 639 unreliable tweets from January 2017 to May 2023. There are no existing dataset that incorporate both text and image AI for misinformation generation. Furthermore, the most datasets rely on website URLs or tweet IDs, which have become partially or fully inaccessible due to Twitter’s API restrictions and ongoing content censorship.
@@ 2.2.1 Methods for textual data
Existing methods typically utilize a latent space to create embedding representations for article content. Embeddings capture contextual information and encode it into a lower-dimensional representation for downstream tasks such as classification and prediction Papanikou et al. (2024). Many approaches leverage this feature using the BERT to achieve exceptional performance in health misinformation classification, either through fine-tuning Kaliyar et al. (2021) or a combination of pre-training and fine-tuning Lee et al. (2019). Other strategies, such as BERTweet Nguyen et al. (2020) and COVID-Twitter-BERT Müller et al. (2023), combine multiple pre-trained models across various contexts. These hybrid models achieve higher performance due to the distinct contextual understanding provided by different models. However, textual models capture only single-modality features, neglecting the visual information often associated with articles. We address this limitation by using SOTA vision-language models to capture both textual and visual context.
@@ 2.2.2 Methods for multimodal data
Previous studies have found that image-containing posts or articles tend to have higher user interaction, spread more quickly, and trigger stronger emotional responses Papanikou et al. (2024). Multimodal misinformation detection requires both image and text encoders to capture textual and visual features. One method Uppada et al. (2023) leverages these encoders to identify fake posts containing visual data, while another approach detects inconsistencies between textual entities and image content for misinformation detection. Limited research has explored multimodal health misinformation detection. DGExplain Shang et al. (2022) is a recent multimodal encoder-decoder architecture for COVID-19 misinformation detection that investigates cross-modal associations between text and images to assess the reliability of news articles. However, these models often require extensive training on benchmark datasets and offer limited interpretability. Our work applies current advanced Vision-Language Models as a baseline for health misinformation detection to evaluate their performance.
@@ 3.1.1 Health Misinformation Collection
To ensure variability, we utilise four recent health-related misinformation datasets Sun et al. (2023); Chen et al. (2021b); Li et al. (2020); Zhou et al. (2020), collected between February 2020 and May 2023, as our primary data sources. Since Twitter (now known as X.com) restricted data access through its application programming interface (API) in 2023, we exclude tweets from these datasets and focus on the available news articles. While Med-MMHL provides raw text and image data, the other datasets only provide URLs as their data source. However, not all URLs are accessible due to news websites consistently removing outdated or misleading news, either voluntarily or in response to complaints. Therefore, we validate the URLs in the database and scrape the text and image metadata from the available URLs and attach the reliable and unreliable labels extracted from the original datasets. We also use the Pillow  [^4] https://python-pillow.org/ library to filter out some images, such as blurry, logo-based, fuzzy, or meaningless images, which are often downloaded during web scraping. See Figure 5 in appendix for examples of the removed images.
@@ 3.1.3 Post-Processing with Human Evaluation
Our two-stage generation process allows AI models to create most of the text and images based on existing data without triggering their internal censorship mechanisms Glukhov et al. (2023). However, some data samples are still not generated, as models may refuse to produce them or output empty strings and black images. Since different models are equipped with different censorship mechanisms, there are variations in behaviour between models. To address this, we developed an algorithm to post-process data alignment between the human-generated context and all five AI-generated contexts (see the Algorithm 1 in appendix for the implementation ). The dataset consists of two sources: human-collected text and images from open-source datasets and AI-generated text-image pairs created by five different models. For consistency, our algorithm will ensured that each AI-generated sample was present across all five models before being included in the final dataset. After data matching, we recruited five evaluators to review the data and further remove low-quality text and image pairs, resulting in 34,746 pieces of health-related information, of which 5,791 were collected from open-source datasets and 28,955 were generated by AI models (see Figure 7 in appendix for examples of reliable and unreliable news generated by humans and AI). The news articles are written in diverse styles by different LLMs while addressing the same topic. While Stable Diffusion 1.5 generates dramatically different images, other models slightly manipulate the original images while maintaining a realistic appearance.
3.2 Data Statistics and Analysis
@@
Table 2: Statistics of generated text and image data from different models. The left table shows text statistics: "Avg Len" is the average word count, and "Avg CSS" is the average Cosine Semantic Similarity. The right table shows image statistics: "FID" is Fréchet Inception Distance and "Avg CIS" is Average Cosine Image Similarity.
Data analysis. To identify the differing data characteristics across the five AI models, we compare their statistics in Table 2. For text models, we calculated the average article length from all five text models. There is a noticeable variation among the models, with ChatGLM-4 tending to generate longer articles, while Mistral-v0.3 is more likely to generate shorter articles. Additionally, we utilise the OpenAI text-embedding-3-small model to generate text embeddings for both human and AI articles, comparing their semantic differences using cosine similarity. For news articles that are too long to fit into the model’s 8,191 context window, we split these long articles into smaller chunks and compute the mean vectors of all chunks to obtain the final text embedding representations for those articles. As indicated on the left side of Table 2, the average cosine semantic similarities across the text models range from 0.76 to 0.78. We plot the Gaussian Kernel Density Estimation (KDE) Zhu et al. (2022) distribution of the semantic similarity in Figure 2, showing that the cosine similarities are concentrated between 0.8 and 0.85 across all models, with slight variance. Notably, a small number of generated articles exhibit low semantic similarity when compared to human generated articles.
For image models, we calculated both the Fréchet Inception Distance (FID) Seitzer (2020) and the average cosine image similarity between AI-generated images and collected images. FID is a common metric used to evaluate the quality of images created from generative models by comparing the distribution of generated images with that of real images. A lower FID score indicates greater similarity between the generated and real images. The average cosine similarity is computed between the vector representations of AI and real images, extracted using ResNet He et al. (2016). The average FID and average cosine image similarities are reported in right side of Table 2, we also plot KDE of the cosine image similarity in Figure 3. We observe that Stable Diffusion 1.5 exhibits the greatest image variance compared to other models, while Stable Diffusion XL and its VAE and PAG variations generate the most similar images. These trends hold for both the average and distribution analyses of the image models. Unlike the text models, no generated images exhibit low similarity when compared to real images.

### v2

@@ Abstract
Infodemics and health misinformation have significant negative impact on individuals and society, exacerbating confusion and increasing hesitancy in adopting recommended health measures. Recent advancements in generative AI, capable of producing realistic, human-like text and images, have significantly accelerated the spread and expanded the reach of health misinformation, resulting in an alarming surge in its dissemination. To combat the infodemics, most existing work has focused on developing misinformation datasets from social media and fact-checking platforms, but has faced limitations in topical coverage, inclusion of AI-generation, and accessibility of raw content. To address these gaps, we present MM-Health, a large scale multimodal misinformation dataset in the health domain consisting of 34,746 news article encompassing both textual and visual information. MM-Health includes human-generated multimodal information (5,776 articles) and AI-generated multimodal information (28,880 articles) from various SOTA generative AI models. Additionally, We benchmarked our dataset against three tasks—reliability checks, originality checks, and fine-grained AI detection—demonstrating that existing SOTA models struggle to accurately distinguish the reliability and origin of information. Our dataset aims to support the development of misinformation detection across various health scenarios, facilitating the detection of human and machine-generated content at multimodal levels [^1] Our code and data is available at: https://github.com/grantzyr/MM-Health-Dataset.
From Generation to Detection: A Multimodal Multi-Task Dataset
for Benchmarking Health Misinformation
Zhihao Zhang1thanks: Equal contribution, Yiran Zhang [^1] footnotemark: 1, Xiyue Zhou2, Liting Huang3, Imran Razzak4, Preslav Nakov4, Usman Naseem1 Macquarie University1, University of Sydney2, UTS3, MBZUAI4 {zhihao.zhang, yiran.zhang, usman.naseem}@mq.edu.au {imran.razzak, preslav.nakov}@mbzuai.ac.ae
@@ 1 Introduction
Health misinformation refers to information that is inaccurate, misleading, or false according to the best available health evidence at the time. Such misinformation tends to spread at unprecedented speed and scale on the World Wide Web and social media Office_of_the_Surgeon_General_OSG2021-op, and has been shown to have significant negative effects on society Borges_do_Nascimento2022-nh; Abbott2021-bg; WANG2019112552; Muhammed_T2022-wu; walter-2020. During crises, such as outbreaks of infectious diseases like COVID-19, the overproduction of health misinformation in both digital and physical environments is defined as an infodemic. As society transitions into a long COVID era Davis2023-qm, the infodemic continues to evolve, leading to a reduction in public trust in health professionals walter-2020. Furthermore, unfiltered exposure to health misinformation can delay or prevent effective disease treatment and even threaten the lives of individuals WANG2019112552. Additionally, the recent advent of generative image models, such as Stable Diffusion stable-diffusion, MidjourneyV5 midjourney, DALL-E2 dall-e, as well as human-level text generators like ChatGPT chat-gpt and LLaMa llama, has amplified both the quantity and spread of misinformation 10.1145/3544548.3581318. The ease with which generative AI models can replicate or manipulate multimodal content, including text and images huang2024ruailargemultimodaldataset, further contributes to the emergence of health misinformation Park2024-ma; ahmad2025vaxguard; shah2024navigating. This presents a significant new challenge in combating the infodemic.
Name Year Human Context Machine Context
Multiple
AI Models
Reliability Originality
General
Avaliable
@@
Recent studies have focused on developing and analyzing datasets and detection methods to combat the infodemic. Common methods for collecting multimodal health misinformation include scraping social media platforms MedHelp; ANTi-Vax; CoAID; ReCOVery; Monant; Med-MMHL; MM-COVID; MMCoVaR, such as Twitter (now known as X), and parsing fact-checking websites MM-COVID; MMCoVaR; Monant; Med-MMHL, such as Snopes and Poynter. However, these datasets exhibit the following notable limitations:
  • Exclusion of AI-generated misinformation: The majority of existing datasets focus solely on human-generated misinformation, ignoring the growing trend of AI-generated misinformation Synthetic; 10.1145/3687028; 10.1145/3581783.3612704. With AI models now capable of generating and manipulating both text and image content LIZLOPEZ2024102103, it is crucial to include AI-generated misinformation to ensure data diversity.
@@
The limitation of the previous health misinformation datasets are summarised in Table 1. Most of the existing methods for automatically detecting health misinformation include content-based CNNs CoAID; ReCOVery; DBLP:conf/icwsm/DaiSW20, attention-based hierarchical encoders Kinkead2020, and graph-based attention methods DETERRENT. These methods have demonstrated competitive performance after being trained on health misinformation benchmarks. However, they require extensive training on benchmark datasets to achieve expected performance, which limits the models’ ability to generalise to new health misinformation. Since health misinformation is constantly evolving Okoro2024ARO, it is important to evaluate generalised models under zero-shot settings. Additionally, these models only accept structured input and provide labelled output without human-readable explanations. This constraint limits the usability of existing detection methods for the general public outside of the research community. To address the aforementioned limitations, we develop MM-Health, a comprehensive multimodal dataset designed for detecting both human and AI generated health misinformation. Additionally, we conduct extensive experiments using state-of-the-art (SOTA) Vision-Language Models (VLLMs) to evaluate their performance in assessing both the reliability and originality of the health information in MM-Health, as well as performing a fine-grained AI detection analysis. Our key contributions are as follows:
  • We create and release a new large-scale multimodal (text and images) dataset called MM-Health, which contains 34,746 news articles. These articles were collected from existing multimodal health datasets and generated using state-of-the-art (SOTA) AI generative models.
@@ 2 Related Work
Web and social media provide valuable sources of information and play important roles in multiple tasks like depression and suicide identification DBLP:journals/tcss/NaseemKKD24, health mention classification DBLP:conf/www/NaseemKKD22 and creditable knowledge management NISAR2019264. However, systematic reviews shows that health-related misinformation on web and social media posts critical threat to the community, which could lead to delay or prevent treatment and even threaten the lives of individuals WANG2019112552.
@@ 2.1 Existing Datasets
MedHelp MedHelp, COAID CoAID, and ANTi-Vax ANTi-Vax are text datasets designed to address online health misinformation. Specifically, MedHelp is collected from an online general health discussion forum. It contains 1,338 non-misinformation comments and 887 misinformation comments annotated by human experts between 2001 and 2013. COAID includes both news articles and Twitter content focused on COVID-19 between December 2019 and September 2020. It labels 204 fake news articles, 3,565 true news articles, 28 fake claims, and 454 true claims from a total of 1,896 news articles, 516 Twitter posts, and 183,569 Twitter engagements. ANTi-Vax is built solely from Twitter content focused on the COVID-19 vaccine between December 2020 and January 2021. It includes a total of 15,073 tweets, 5,751 of which are misinformation and 9,322 are general vaccine-related tweets.
Figure 1: Data collection process for $M$-Health includes: 1) utilising multiple existing open-source health misinformation datasets as data sources, 2) validating the available data samples and collecting human-generated multimodal data from the provided URLs, and 3) implementing generative AI models to collect AI-generated replicated multimodal data. To ensure data quality, both human and AI generated content are evaluated by five human evaluators proficient in English.
MM-COVID MM-COVID, ReCOVery ReCOVery, Monant Monant, MMCoVaR MMCoVaR, and Med-MMHL Med-MMHL are multimodal datasets for health-related misinformation detection. MM-COVID contains 3,981 fake articles and 7,584 real articles related to COVID-19 which collected from Feb 2020 to Jul 2020. ReCOVery includes both news articles and tweets mentioning COVID-19, labelled by NewsGuard  [^3] https://www.newsguardtech.com/ and Media Bias/Fact Check  [^4] https://mediabiasfactcheck.com/. It comprises 118,339 reliable articles and 28,274 unreliable articles sourced between January 2020 and May 2020. MMCoVaR focuses on vaccine-related misinformation between February 2020 and March 2021, featuring 958 unreliable and 1,635 reliable news articles, plus 24,184 tweets categorized as reliable, unreliable, or inconclusive. Med-MMHL covers multiple diseases, integrating LLM-generated fake news, with 1,979 real and 1,604 fake articles, alongside 1,334 reliable and 639 unreliable tweets from January 2017 to May 2023. There are no existing dataset that incorporate both text and image AI for misinformation generation. Furthermore, the most datasets rely on website URLs or tweet IDs, which have become partially or fully inaccessible due to Twitter’s API restrictions and ongoing content censorship.
@@ 2.2.1 Methods for Textual Data
Existing methods typically utilize a latent space to create embedding representations for article content. Embeddings capture contextual information and encode it into a lower-dimensional representation for downstream tasks such as classification and prediction papanikou2024healthmisinformationsocialnetworks. Many approaches leverage this feature using the BERT to achieve exceptional performance in health misinformation classification, either through fine-tuning FakeBERT or a combination of pre-training and fine-tuning BioBERT. Other strategies, such as BERTweet bertweet and COVID-Twitter-BERT COVID-Twitter-BERT, combine multiple pre-trained models across various contexts. These hybrid models achieve higher performance due to the distinct contextual understanding provided by different models. However, textual models capture only single-modality features, neglecting the visual information often associated with articles. We address this limitation by using SOTA vision-language models to capture both textual and visual context.
@@ 2.2.2 Methods for Multimodal Data
Previous studies have found that image-containing posts or articles tend to have higher user interaction, spread more quickly, and trigger stronger emotional responses papanikou2024healthmisinformationsocialnetworks. Multimodal misinformation detection requires both image and text encoders to capture textual and visual features. One method Uppada2023 leverages these encoders to identify fake posts containing visual data, while another approach detects inconsistencies between textual entities and image content for misinformation detection. Limited research has explored multimodal health misinformation detection. DGExplain 10.1145/3485447.3512257 is a recent multimodal encoder-decoder architecture for COVID-19 misinformation detection that investigates cross-modal associations between text and images to assess the reliability of news articles. However, these models often require extensive training on benchmark datasets and offer limited interpretability. Our work applies current advanced Vision-Language Models as a baseline for health misinformation detection to evaluate their performance.
@@ 3.1.1 Health Misinformation Collection
To ensure variability, we utilise four recent health-related misinformation datasets Med-MMHL; MMCoVaR; MM-COVID; ReCOVery, collected between February 2020 and May 2023, as our primary data sources. Since Twitter (now known as X.com) restricted data access through its application programming interface (API) in 2023, we exclude tweets from these datasets and focus on the available news articles. While Med-MMHL provides raw text and image data, the other datasets only provide URLs as their data source. However, not all URLs are accessible due to news websites consistently removing outdated or misleading news, either voluntarily or in response to complaints. Therefore, we validate the URLs in the database and scrape the text and image metadata from the available URLs and attach the reliable and unreliable labels extracted from the original datasets. We also use the Pillow [^5] https://python-pillow.org/ library to filter out some images, such as blurry, logo-based, or meaningless images, which are often downloaded during web scraping. Examples of removed images are in Appendix (Figure 5).
@@ 3.1.3 Post-Processing with Human Evaluation
Our two-stage generation process allows AI models to create most of the text and images based on existing data without triggering their internal censorship mechanisms glukhov2023llmcensorshipmachinelearning. However, some data samples are still not generated, as models may refuse to produce them or output empty strings and black images. Since different models are equipped with different censorship mechanisms, there are variations in behaviour between models. To address this, we developed an algorithm to post-process data alignment between the human-generated context and all five AI-generated contexts(see the Algorithm 1 in appendix for the implementation ). The dataset consists of two sources: human-collected text and images from open-source datasets and AI-generated text-image pairs created by five different models. For consistency, our algorithm will ensured that each AI-generated sample was present across all five models before being included in the final dataset. After data matching, human evaluation was conducted by five independent annotators. These annotators were students ranging from junior to senior levels, including both male and female participants. All had backgrounds in AI literature and a foundational understanding of health-related topics, which helped ensure informed and thorough evaluation. Before the annotation task, all participants underwent thorough training to familiarize them with the evaluation criteria, common failure modes (e.g., logo artifacts, mismatched captions), and the overall goal of the task. This training ensured consistency and reliability in their assessments. The annotators were compensated fairly according to the university’s standard participant payment guidelines. Inter-annotator agreement was measured and found to be high, with a score above 80%, indicating strong consistency across evaluators. While domain-specific expertise was not strictly required, the annotators’ familiarity with AI and health contexts contributed to a more nuanced understanding of the generated image–text pairs. This results in 34,746 pieces of health-related information, of which 5,791 were collected from open-source datasets and 28,955 were generated by AI models Examples of reliable and unreliable news generated by humans and AI are in Figure 7 in the appendix. The news articles are written in diverse styles by different LLMs while addressing the same topic. While Stable Diffusion 1.5 generates dramatically different images, other models slightly manipulate the original images while maintaining a realistic appearance.
@@
Table 2: Statistics of generated text and image data from different models. The left table shows text statistics: "Avg Len" is average word count, and "Avg CSS" is average Cosine Semantic Similarity. The right table shows image statistics: "FID" is Fréchet Inception Distance and "Avg CIS" is Average Cosine Image Similarity.
3.2 Data Statistics and Analysis
Data analysis. To identify the differing data characteristics across the five AI models, we compare their statistics in Table 2. For text models, we calculated the average article length from all five text models. There is a noticeable variation among the models, with ChatGLM-4 tending to generate longer articles, while Mistral-v0.3 is more likely to generate shorter articles. Additionally, we utilise the OpenAI text-embedding-3-small model to generate text embeddings for both human and AI articles, comparing their semantic differences using cosine similarity. For news articles that are too long to fit into the model’s 8,191 context window, we split these long articles into smaller chunks and compute the mean vectors of all chunks to obtain the final text embedding representations for those articles. As indicated on the left side of Table 2, the average cosine semantic similarities across the text models range from 0.76 to 0.78. We plot the Gaussian Kernel Density Estimation (KDE) DBLP:journals/isci/ZhuLHK22 distribution of the semantic similarity in Figure 2, showing that the cosine similarities are concentrated between 0.8 and 0.85 across all models, with slight variance. Notably, a small number of generated articles exhibit low semantic similarity when compared to human generated articles.
For image models, we calculated both the Fréchet Inception Distance (FID) Seitzer2020FID and the average cosine image similarity between AI-generated images and collected images. FID is a common metric used to evaluate the quality of images created from generative models by comparing the distribution of generated images with that of real images. A lower FID score indicates greater similarity between the generated and real images. The average cosine similarity is computed between the vector representations of AI and real images, extracted using ResNet resnet. The average FID and average cosine image similarities are reported in right side of Table 2, we also plot KDE of the cosine image similarity in Figure 3. We observe that Stable Diffusion 1.5 exhibits the greatest image variance compared to other models, while Stable Diffusion XL and its VAE and PAG variations generate the most similar images. These trends hold for both the average and distribution analyses of the image models. Unlike the text models, no generated images exhibit low similarity when compared to real images.